# A Survey on Safe Multi-Modal Learning Systems

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 4.3. Benchmark
In the domain of failure detection, datasets typically consist of normal data without explicitly labeled anomaly subsets. For failure detection in MAD, the choice of datasets varies, reflecting diverse areas of study encompassing medical imaging to image-text corpora. Conversely, in AD scenarios, datasets are specifically annotated to highlight anomalies. When evaluating Vision-Language MAD, commonly used benchmark datasets include MVTec-AD (Bergmann et al., 2019), VisA (Zou et al., 2022), KSDD2 (Božič et al., 2021), MPDD (Jezek et al., 2021), MTD (Huang et al., 2020), BTAD (Mishra et al., 2021), etc. These industrial inspection datasets are annotated at the pixel level and cover a wide range of object subsets. Datasets for Video Surveillance MAD include ShanghaiTech (Liu et al., 2018), UCF-Crime (Sultani et al., 2018) and XD-Violence (Wu et al., 2020), an audiovisual abnormal dataset. Performance evaluation is primarily based on metrics such as the area under the ROC curve, area under the precision-recall curve, and F1 score. Other metrics include Per-Region Overlap scores, which assess anomaly segmentation performance, and Average Precision.
@@ 8.2. Preserving Privacy in MMLS
Relevant research primarily focuses on differential privacy, machine unlearning, and federated learning.
@@ 8.3. Benchmark
Current memorization measurements lack standardization and are primarily restricted to VLMs. Common metrics include Frechet Inception Distance score (Somepalli et al., 2023) and Memorization score (Zhang et al., 2023a). Machine unlearning evaluation parallels that in unimodal settings, including attack-based methods like Membership Inference Robustness, and accuracy-based metrics like Mean Squared Error, alongside others like Activation Distance and Retraining Time (Nguyen et al., 2022). Benchmarking multimodal federated learning is still in its early stages. FedMultimodal is the first systematic benchmark (Feng et al., 2023), offering 10 representative datasets from 5 key MMLS applications (emotion recognition, action recognition, human activity recognition, healthcare, and social media) and providing a framework covering the pipeline of multimodal federated learning for effective assessment.
@@ 9. Conclusion and Open Challenges
This paper presents the first comprehensive survey on safe MMLS and introduces a systematic taxonomy centering on four critical dimensions: robustness, alignment, monitoring, and controllability. Following this taxonomy, we review current research, summarizing existing progress and challenges. We hope that this survey encourages additional advancements in the realm of safe MMLS. Below, we highlight several challenges in MMLS safety, aiming to offer insights that can serve as valuable guidance for future research.
@@
Differentially Private Training. Differential privacy techniques demonstrate their effectiveness in safeguarding privacy in unimodal systems. Yet, adapting these techniques for MMLS faces challenges and is in its infancy. Some strategies to enhance differential privacy are to apply modality-specific noise addition and coordinate across modalities by considering the interdependence between modalities and developing mechanisms that preserve privacy across the multimodal model without sacrificing utility.
@@
Advancing Datasets for MMLS Safety. The challenges associated with multimodal datasets encompass various aspects, such as the lack of available open-source large models capable of effectively handling multimodal inputs. Current models produce content that lacks interpretability, impeding understanding and trust. Ensuring the quality control of generated data poses difficulties, demanding significant resources for collection efforts. Research in MMLS tends to be limited to specific use cases with an uneven distribution of datasets complicating the generalization of models, particularly in terms of fairness. This mirrors the challenges faced by unimodal models. To address these issues, further research and refinement of public datasets are essential.

### v4

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; August 25–29, 2024; Barcelona, Spain.booktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spainisbn: 979-8-4007-0490-1/24/08doi: 10.1145/3637528.3671462ccs: Computing methodologies Artificial intelligence
@@ 4.3. Benchmark
In failure detection, datasets typically consist of normal data without explicitly labeled anomaly subsets. For failure detection in MAD, datasets vary widely, covering areas from medical imaging to image-text corpora. Conversely, in AD scenarios, datasets are specifically annotated to highlight anomalies. When evaluating Vision-Language MAD, commonly used benchmark datasets include MVTec-AD (Bergmann et al., 2019), VisA (Zou et al., 2022), KSDD2 (Božič et al., 2021), MPDD (Jezek et al., 2021), MTD (Huang et al., 2020), etc. These industrial inspection datasets are annotated at the pixel level and cover a wide range of object subsets. Datasets for Video Surveillance MAD include ShanghaiTech (Liu et al., 2018), UCF-Crime (Sultani et al., 2018) and XD-Violence (Wu et al., 2020), an audiovisual abnormal dataset. Performance evaluation is primarily based on metrics such as the area under the ROC curve, area under the precision-recall curve, and F1 score. Other metrics include Per-Region Overlap scores, which assess anomaly segmentation performance, and Average Precision.
@@ 8.3. Benchmark
Current memorization measurements lack standardization and are primarily restricted to VLMs. Common metrics include Frechet Inception Distance score (Somepalli et al., 2023) and Memorization score (Zhang et al., 2023a). Machine unlearning evaluation parallels that in unimodal settings, including attack-based methods like Membership Inference Robustness, and accuracy-based metrics like Mean Squared Error, alongside others like Activation Distance and Retraining Time (Nguyen et al., 2022). Benchmarking multimodal federated learning is still in its early stages. FedMultimodal is the first systematic benchmark (Feng et al., 2023), offering 10 representative datasets from 5 key MMLS applications and providing a pipeline for effective multimodal federated learning assessment.
@@ 9. Conclusion and Open Challenges
This paper presents the first comprehensive survey on safe MMLS, introducing a systematic taxonomy centering on four critical dimensions: robustness, alignment, monitoring, and controllability. Following this taxonomy, we review and summarize current research and progress. We hope that this survey offer insights for future research. Below, we highlight several challenges in MMLS safety.
@@
Differentially Private Training. Differential privacy techniques demonstrate their effectiveness in safeguarding privacy in unimodal systems. Yet, adapting these techniques for MMLS faces challenges and is in its infancy. Some strategies to enhance differential privacy are to apply modality-specific noise addition and coordinate across modalities by considering the interdependence between modalities.
@@
Advancing Datasets for MMLS Safety. Challenges with multimodal datasets include the scarcity of open-source large models capable of effectively handling multimodal inputs, limited interpretability of existing models that hinders understanding and trust, and the difficulty of quality control and data collection. In addition, research in MMLS is often restricted to specific use cases, and uneven dataset distribution hinders model generalization. These issues require more efforts in data collection and refinement.
@@
Acknowledgements
This material is based upon work supported by the National Science Foundation (NSF) Grant #2312862, #2406648, #2406647 and a Cisco gift grant.